OpenStudy (anonymous):

The amount of material to cover a spherical ball with a diameter of 2.9 inches is 26.4 inches2. If the diameter of the ball is tripled, how much material will be required to cover the new surface of the ball?

jhonyy9 (jhonyy9):

diameter 2,9 inches --------------- 26,4 inches squared material diameter tripled 3*2,9=8,7 inches ----- x -------------------------------------------------------- x = 8,7 * 26,4 / 2,9 = 79,2 inches squared material will be required to cover the new surface og the ball